# Preprint submitted to Computers & Geosciences

### v4

@@ 1 Introduction
A variety of low-dimensional parameterization methods have been proposed over the years. Previous treatments include, e.g., the use of discrete cosine transform (Jafarpour and McLaughlin, 2009) and principal component analysis (PCA) and its variants, such as K-PCA and O-PCA (Sarma et al., 2008; Vo and Durlofsky, 2015). Although useful in many settings, these methods have limited accuracy in maintaining complex geological features in non-Gaussian systems (Chan and Elsheikh, 2019), particularly when there is a limited amount of hard data for conditioning.
@@
Recently, deep learning models have emerged as powerful alternatives for geological parameterization. Some of the methods relevant to this study include variatonal autoencoders, VAEs (Canchumuni et al., 2019; Mo et al., 2020; Liu and Grana, 2020), in which a neural network learns the mapping from the geomodel data distribution to a latent space with a known distribution (and vice versa), CNN-PCA approaches (Liu and Durlofsky, 2021; Tang et al., 2021), which involve a convolutional neural network post-processing of a PCA representation, and generative adversarial networks, or GANs (Laloy et al., 2018; Zhang et al., 2019, 2021; Chan and Elsheikh, 2019; Song et al., 2021b, 2023; Hu et al., 2023; Feng et al., 2024). GANs have received substantial attention and they perform well in many cases. However, the adversarial training they require is typically time-consuming and may be unstable, which necessitates careful hyperparameter tuning. CNN-PCA methods currently treat only a single geological scenario at a time, and their performance depends on the amount of conditioning data (as is also the case for other PCA-based methods). Finally, while VAEs alone are generally superior to standard PCA for non-Gaussian models, they can produce blurry output of lower quality than GANs, or they may display unrealistic geometries (Canchumuni et al., 2019).
@@
In the last few years, generative diffusion models, introduced by Ho et al. (2020), have become the state-of-the-art for image generation tasks, outperforming the established GANs in several benchmark studies (Dhariwal and Nichol, 2021). Diffusion methods, which do not require adversarial training, have found successful application in many fields of science and engineering that involve structured 2D or 3D data. This includes image and video generation, molecule and material design, super-resolution, and time-series forecasting (Yang et al., 2023). These models are especially suitable for conditioned generative tasks, where conditioning can be in the form of scalars, text or images. These properties make diffusion models ideal candidates for geomodeling and parameterization. Diffusion models entail two discrete-step processes: noising (forward process) and denoising (reverse process). Noising progressively adds noise to the data, and the model is trained to learn the reverse process. In this way, diffusion models can generate new samples from a random noise input. There have been only a few published studies that have used these methods for subsurface flow applications. Mosser (2023), for example, applied diffusion models for geostatistical simulation of 2D fluvial channel and river datasets. However, the latent space in these treatments had the same dimensionality as the geomodel space, limiting their application to small models.
@@
To address this issue, Lee et al. (2023) proposed the use of latent diffusion models, or LDMs, developed by Rombach et al. (2022). LDMs operate in the same way as standard diffusion models, but they use a low-dimensional latent variable in conjunction with an encoder-decoder network architecture. Lee et al. (2023) applied LDMs for the conditional generation of multifacies 2D geomodels. This method, however, relies on a random sequence of sampling steps (known as denoising diffusion probabilistic model, DDPM). In addition, a large number of steps is required to achieve high geomodel quality, resulting in long generation times. These features may limit the applicability of this approach for history matching problems. Song et al. (2021a) introduced denoising diffusion implicit models (DDIMs), which represent a compromise between computational time and accuracy. DDIMs enable faster, deterministic sampling, with added latent space interpolation capabilities and image quality that is comparable to DDPMs (Song et al., 2021a). To our knowledge, DDIM sampling in combination with LDMs has not been considered for geomodel parameterization or data assimilation.
@@
In this work, we extend and apply LDMs to parameterize 2D multifacies (channel-levee-mud) geomodels. In contrast to previous implementations, our workflow entails fast deterministic sampling and smooth/stable parameterization, both of which are important for data assimilation. Training samples are obtained using object-based modeling in Petrel. Metrics involving spatial statistics and flow quantities are used to assess the quality of the LDM geomodels relative to (reference) object-based geomodels. History matching, specifically ESMDA (Emerick and Reynolds, 2013) applied to the latent variables, is performed with the LDM geomodels for two-phase (oil-water) problems.
@@ 2 Diffusion-model-based parameterization
In this section, we first review the DDPM (Ho et al., 2020) diffusion generative process. Next, we describe the DDIM (Song et al., 2021a) sampling procedure and LDM parameterization (Rombach et al., 2022). Finally, we present the specific approach and model architecture adopted in this work.
@@ 2.1 Denoising diffusion probabilistic models (DDPMs)
Consider a set of conditioned geomodels, represented on a discrete grid comprised of $N_{x}\times N_{y}=N_{c}$ cells. We denote a sample (realization) as $\mathbf{m}_{0}\in\mathbb{R}^{N_{x}\times N_{y}}$, where the 00 subscript refers to the clean (without noise) geomodel. The distribution characterizing the geomodel, $q(\mathbf{m}_{0})$, is unknown. Starting from $\mathbf{m}_{0}$, the noising (forward) process $q$ over $T$ discrete steps is modeled as a Markov chain of discrete Gaussian transition steps (Ho et al., 2020):
@@
[…]
@@
The denoising (inverse) process $p_{\theta}$ is learned by a neural network, where subscript $\theta$ denotes the network parameters. Starting from noise $p_{\theta}(\mathbf{m}_{T})$, this can be expressed as (Ho et al., 2020):
@@
Here $\bm{\mu}_{\theta}$ and $\bm{\Sigma}_{\theta}$ are the learnable parameters of the inverse Gaussian steps. In Ho et al. (2020), the $\bm{\Sigma}_{\theta}$ parameters are taken as $\sigma_{t}^{2}\bf{I}$, with $\sigma_{t}^{2}=\beta_{t}$, such that they are constant (not learned) and defined by the scheduler. The means ($\bm{\mu}_{\theta}$), by contrast, are learned and parameterized as:
@@
In this way, when the network is provided with a clean model $\mathbf{m}_{0}$ that has been noised until step $t$ (which gives $\mathbf{m}_{t}$), the diffusion model is trained to predict the noise $\bm{\epsilon}_{\theta}$. This noise will then be subtracted during denoising. Training is performed by minimizing the loss function between true noise (added by the scheduler and therefore known) and predicted noise (Ho et al., 2020):
$L(\theta)=\mathbb{E}_{\mathbf{m}_{t},\bm{\epsilon}}\left[\Big{|}\Big{|}\bm{\epsilon}-\bm{\epsilon}_{\theta}(\mathbf{m}_{t},t)\Big{|}\Big{|}_{2}^{2}\right]=\mathbb{E}_{t,\mathbf{m}_{0},\bm{\epsilon}}\left[\Big{|}\Big{|}\bm{\epsilon}-\bm{\epsilon}_{\theta}\left(\sqrt{\bar{\alpha}_{t}}\mathbf{m}_{0}+\sqrt{1-\bar{\alpha}_{t}}\bm{\epsilon},t\right)\Big{|}\Big{|}_{2}^{2}\right].$ (5)
@@ 2.2 Denoising diffusion implicit models (DDIMs)
In the DDPM procedure described above, the number of steps $T$ required to achieve high-quality geomodels can be $O(10^{3})$. In addition, Eq. 6 contains the random component $\mathbf{z}$, which renders the generation process nondeterministic, even when starting with the same initial noise $\mathbf{m}_{T}$. This may be a limitation for data assimilation problems, where updates to the input variables ($\mathbf{m}_{T}$ in this case) should affect the generated output ($\mathbf{m}_{0}$) in a predictable manner. To overcome this limitation, Song et al. (2021a) introduced denoising diffusion implicit models (DDIMs). They suggested reparameterizing the reverse diffusion in Eq. 6 from a Markov chain to a non-Markovian expression for a single step $\mathbf{m}_{t-1}$:
@@
where $\bar{\alpha}_{t}=\prod_{s=1}^{t}\alpha_{s}$, and similarly for $\bar{\alpha}_{t-1}$. DDIMs maintain image quality that is comparable to that of DDPMs (Song et al., 2021a), with the advantages of faster sampling (with as few as 20-50 steps), higher consistency in the generative process ($\mathbf{m}_{0}$ depends only on the initial state $\mathbf{m}_{T}$), and interpolation capabilities. In addition, because the training objective with the DDIM is the same as that for the DDPM, a DDIM can be used for inference on a DDPM-trained diffusion model. More details can be found in Song et al. (2021a).
@@ 2.3 Latent diffusion models (LDMs)
In the diffusion models described thus far, the latent space has the same dimensionality as the geomodel space, i.e., the parameterization involves a mapping, but no dimension reduction. For data assimilation applications, it is beneficial to reduce the number of geomodel parameters that must be determined, and many previous parameterization methods provide this capability. The LDMs introduced by Rombach et al. (2022) are very useful in this regard. They have the same structure as standard diffusion models, but the noising and denoising operations are performed on a lower-dimension latent variable, $\bm{\xi}\in\mathbb{R}^{n_{x}\times n_{y}}$, where $n_{x}=N_{x}/f$ and $n_{y}=N_{y}/f$. Here $f$ is referred to as the downsampling ratio. An encoder $\mathcal{E}(\mathbf{m}):\mathbb{R}^{N_{x}\times N_{y}}\rightarrow\mathbb{R}^{n_{x}\times n_{y}}$ and a decoder $\mathcal{D}(\bm{\xi}):\mathbb{R}^{n_{x}\times n_{y}}\rightarrow\mathbb{R}^{N_{x}\times N_{y}}$, as shown in Figure 2a, perform these mappings. A schematic of the full LDM is shown in Figure 2b.
@@
Figure 2: Schematic representation of encoder-decoder and full LDM architecture (inspired by Rombach et al. (2022)).
The LDM used in this study consists of a variational autoencoder (VAE) and a U-net (Ronneberger et al., 2015). The VAE performs the dimension reduction from the geomodel space to the latent space and vice-versa (Figure 2(a)). The U-net is used for the reverse diffusion process (Figure 2(b)). The VAE consists of an encoder and a decoder. The encoder is comprised of three downsampling convolutional layers, alternated with four residual blocks (group normalization – convolution – group normalization – convolution), such that the total downsampling ratio $f=8$ and the latent space $\bm{\xi}$ is of dimensions $N_{x}/8\times N_{y}/8$. Note that, in contrast to most VAE implementations, which rely on a 1D latent space, LDMs use a 2D latent space. The decoder mirrors the encoder, and thus has four residual blocks and three upsampling convolutional layers. The U-net consists of two downsampling residual blocks, one middle block with an attention mechanism (Bahdanau et al., 2015), and two upsampling residual blocks. The overall architecture follows the implementation of the original LDM in Rombach et al. (2022). Our implementation is developed using the Python libraries diffusers and monai (see the ”Code availability” section for additional details). The detailed VAE and U-net architectures are shown in Figure 3.
@@ 2.4 LDM loss function and training
$L_{recon}=||\mathbf{m}_{0}-\hat{\mathbf{m}}_{0}||_{2}^{2}.$ (8)
The latent variable $\bm{\xi}$ is intended to be standard normal (Rombach et al., 2022), and the K-L divergence quantifies the difference between the distribution provided by the encoder and $\mathcal{N}(\mathbf{0},\mathbf{I})$, i.e.,
@@
$L_{LDM}(\theta)=\mathbb{E}_{t,\mathcal{E}(\mathbf{m}_{0}),\bm{\epsilon}}\left[\Big{|}\Big{|}\bm{\epsilon}-\bm{\epsilon}_{\theta}\left(\sqrt{\bar{\alpha}_{t}}\bm{\xi}_{0}+\sqrt{1-\bar{\alpha}_{t}}\bm{\epsilon},t\right)\Big{|}\Big{|}^{2}_{2}\right].$ (12)
In practice, during training, all losses are computed in batches of $N_{b}$ models. Pseudo-code for the U-net training is provided in Algorithm 1 (modified from Ho et al. (2020)). For inference (generation), denoising starts from a lower dimensionality latent vector $\bm{\xi}_{T}$. The denoising in Eq. 7 is applied on $\bm{\xi}$. At the final step it is transformed by the decoder into a full-size geomodel $\mathbf{m}_{DM}$. This procedure is summarized in Algorithm 2 (modified from Ho et al. (2020)).
@@
The dataset used in this study consists of 4000 conditional realizations, of dimensions $64\times 64$ ($N_{x}\times N_{y}=N_{c}=4096$), of three-facies geomodels. The facies correspond to high-permeability channel sand, low-permeability background mud, and intermediate-permeability levee features. The training realizations are generated with Petrel software (Schlumberger, 2019) using object-based modeling. These realizations are of similar style to those used by Song et al. (2021b). A 70-20-10% split for training, validation, and testing is applied. There are five hard-data locations, where the models are conditioned to channel facies. Five training realizations are shown in Figure 4.
The total training time is about 2 h (of which 0.5 h is for VAE training and 1.5 h is for U-net training) on one Nvidia A100 GPU. We use the Adam optimizer (Kingma and Ba, 2015), with a learning rate of 1e-4 and a batch size of 16. The scheduler for the noising process is linear, with $\beta_{1}=0.0001$ and $\beta_{T}=0.02$, as in Ho et al. (2020). A total of 100 DDIM steps are applied. The weights $\lambda_{KL}$ and $\lambda_{h}$ are set to $10^{-6}$ (as in Rombach et al. (2022)) and $10$, respectively.

### v5

# Latent diffusion models for parameterization and data assimilation of facies-based geomodels

@@ 1 Introduction
A variety of low-dimensional parameterization methods have been proposed over the years. Previous treatments include, e.g., the use of discrete cosine transform (Jafarpour) and principal component analysis (PCA) and its variants, such as K-PCA and O-PCA (Sarma2008; Vo2015). Although useful in many settings, these methods have limited accuracy in maintaining complex geological features in non-Gaussian systems (Chan_2019), particularly when there is a limited amount of hard data for conditioning.
@@
Recently, deep learning models have emerged as powerful alternatives for geological parameterization. Some of the methods relevant to this study include variatonal autoencoders, VAEs (Canchumuni_2019; mo_caae; grana_co2), in which a neural network learns the mapping from the geomodel data distribution to a latent space with a known distribution (and vice versa), CNN-PCA approaches (LIU2021104676; Tang_2021), which involve a convolutional neural network post-processing of a PCA representation, and generative adversarial networks, or GANs (Laloy_2018; chan2019parametrization; zhang; Chan_2019; Song_Mukerji_Hou_2021; SONG2023129493; HU2023105290; Feng2024). GANs have received substantial attention and they perform well in many cases. However, the adversarial training they require is typically time-consuming and may be unstable, which necessitates careful hyperparameter tuning. CNN-PCA methods currently treat only a single geological scenario at a time, and their performance depends on the amount of conditioning data (as is also the case for other PCA-based methods). Finally, while VAEs alone are generally superior to standard PCA for non-Gaussian models, they can produce blurry output of lower quality than GANs, or they may display unrealistic geometries (Canchumuni_2019).
@@
In the last few years, generative diffusion models, introduced by ho2020denoising, have become the state-of-the-art for image generation tasks, outperforming the established GANs in several benchmark studies (dhariwal2021diffusion). Diffusion methods, which do not require adversarial training, have found successful application in many fields of science and engineering that involve structured 2D or 3D data. This includes image and video generation, molecule and material design, super-resolution, and time-series forecasting (review_dms). These models are especially suitable for conditioned generative tasks, where conditioning can be in the form of scalars, text or images. These properties make diffusion models ideal candidates for geomodeling and parameterization. Diffusion models entail two discrete-step processes: noising (forward process) and denoising (reverse process). Noising progressively adds noise to the data, and the model is trained to learn the reverse process. In this way, diffusion models can generate new samples from a random noise input. There have been only a few published studies that have used these methods for subsurface flow applications. mosser, for example, applied diffusion models for geostatistical simulation of 2D fluvial channel and river datasets. However, the latent space in these treatments had the same dimensionality as the geomodel space, limiting their application to small models.
@@
To address this issue, lee2023latent proposed the use of latent diffusion models, or LDMs, developed by rombach2022highresolution. LDMs operate in the same way as standard diffusion models, but they use a low-dimensional latent variable in conjunction with an encoder-decoder network architecture. lee2023latent applied LDMs for the conditional generation of multifacies 2D geomodels. This method, however, relies on a random sequence of sampling steps (known as denoising diffusion probabilistic model, DDPM). In addition, a large number of steps is required to achieve high geomodel quality, resulting in long generation times. These features may limit the applicability of this approach for history matching problems. song2021denoising introduced denoising diffusion implicit models (DDIMs), which represent a compromise between computational time and accuracy. DDIMs enable faster, deterministic sampling, with added latent space interpolation capabilities and image quality that is comparable to DDPMs (song2021denoising). To our knowledge, DDIM sampling in combination with LDMs has not been considered for geomodel parameterization or data assimilation.
@@
In this work, we extend and apply LDMs to parameterize 2D multifacies (channel-levee-mud) geomodels. In contrast to previous implementations, our workflow entails fast deterministic sampling and smooth/stable parameterization, both of which are important for data assimilation. Training samples are obtained using object-based modeling in Petrel. Metrics involving spatial statistics and flow quantities are used to assess the quality of the LDM geomodels relative to (reference) object-based geomodels. History matching, specifically ESMDA (EMERICK20133) applied to the latent variables, is performed with the LDM geomodels for two-phase (oil-water) problems.
@@ 2 Diffusion-model-based parameterization
In this section, we first review the DDPM (ho2020denoising) diffusion generative process. Next, we describe the DDIM (song2021denoising) sampling procedure and LDM parameterization (rombach2022highresolution). Finally, we present the specific approach and model architecture adopted in this work.
@@ 2.1 Denoising diffusion probabilistic models (DDPMs)
Consider a set of conditioned geomodels, represented on a discrete grid comprised of $N_{x}\times N_{y}=N_{c}$ cells. We denote a sample (realization) as $\mathbf{m}_{0}\in\mathbb{R}^{N_{x}\times N_{y}}$, where the 0 subscript refers to the clean (without noise) geomodel. The distribution characterizing the geomodel, $q(\mathbf{m}_{0})$, is unknown. Starting from $\mathbf{m}_{0}$, the noising (forward) process $q$ over $T$ discrete steps is modeled as a Markov chain of discrete Gaussian transition steps (ho2020denoising):
@@
$q(\mathbf{m}_{t}|\mathbf{m}_{0})=\int q(\mathbf{m}_{1:t}|\mathbf{m}_{0})d\mathbf{m}_{1:t-1}=\mathcal{N}(\mathbf{m}_{t};\sqrt{\bar{\alpha}_{t}}\mathbf{m}_{0},(1-\bar{\alpha}_{t})\mathbf{I}),$ (2)
@@
The denoising (inverse) process $p_{\theta}$ is learned by a neural network, where subscript $\theta$ denotes the network parameters. Starting from noise $p_{\theta}(\mathbf{m}_{T})$, this can be expressed as (ho2020denoising):
@@
Here $\bm{\mu}_{\theta}$ and $\bm{\Sigma}_{\theta}$ are the learnable parameters of the inverse Gaussian steps. In ho2020denoising, the $\bm{\Sigma}_{\theta}$ parameters are taken as $\sigma_{t}^{2}\bf{I}$, with $\sigma_{t}^{2}=\beta_{t}$, such that they are constant (not learned) and defined by the scheduler. The means ($\bm{\mu}_{\theta}$), by contrast, are learned and parameterized as:
@@
In this way, when the network is provided with a clean model $\mathbf{m}_{0}$ that has been noised until step $t$ (which gives $\mathbf{m}_{t}$), the diffusion model is trained to predict the noise $\bm{\epsilon}_{\theta}$. This noise will then be subtracted during denoising. Training is performed by minimizing the loss function between true noise (added by the scheduler and therefore known) and predicted noise (ho2020denoising):
[…]
@@ 2.2 Denoising diffusion implicit models (DDIMs)
In the DDPM procedure described above, the number of steps $T$ required to achieve high-quality geomodels can be $O(10^{3})$. In addition, Eq. 6 contains the random component $\mathbf{z}$, which renders the generation process nondeterministic, even when starting with the same initial noise $\mathbf{m}_{T}$. This may be a limitation for data assimilation problems, where updates to the input variables ($\mathbf{m}_{T}$ in this case) should affect the generated output ($\mathbf{m}_{0}$) in a predictable manner. To overcome this limitation, song2021denoising introduced denoising diffusion implicit models (DDIMs). They suggested reparameterizing the reverse diffusion in Eq. 6 from a Markov chain to a non-Markovian expression for a single step $\mathbf{m}_{t-1}$:
@@
where $\bar{\alpha}_{t}=\prod_{s=1}^{t}\alpha_{s}$, and similarly for $\bar{\alpha}_{t-1}$. DDIMs maintain image quality that is comparable to that of DDPMs (song2021denoising), with the advantages of faster sampling (with as few as 20-50 steps), higher consistency in the generative process ($\mathbf{m}_{0}$ depends only on the initial state $\mathbf{m}_{T}$), and interpolation capabilities. In addition, because the training objective with the DDIM is the same as that for the DDPM, a DDIM can be used for inference on a DDPM-trained diffusion model. More details can be found in song2021denoising.
@@ 2.3 Latent diffusion models (LDMs)
In the diffusion models described thus far, the latent space has the same dimensionality as the geomodel space, i.e., the parameterization involves a mapping, but no dimension reduction. For data assimilation applications, it is beneficial to reduce the number of geomodel parameters that must be determined, and many previous parameterization methods provide this capability. The LDMs introduced by rombach2022highresolution are very useful in this regard. They have the same structure as standard diffusion models, but the noising and denoising operations are performed on a lower-dimension latent variable, $\bm{\xi}\in\mathbb{R}^{n_{x}\times n_{y}}$, where $n_{x}=N_{x}/f$ and $n_{y}=N_{y}/f$. Here $f$ is referred to as the downsampling ratio. An encoder $\mathcal{E}(\mathbf{m}):\mathbb{R}^{N_{x}\times N_{y}}\rightarrow\mathbb{R}^{n_{x}\times n_{y}}$ and a decoder $\mathcal{D}(\bm{\xi}):\mathbb{R}^{n_{x}\times n_{y}}\rightarrow\mathbb{R}^{N_{x}\times N_{y}}$, as shown in Figure 2a, perform these mappings. A schematic of the full LDM is shown in Figure 2b.
@@
Figure 2: Schematic representation of encoder-decoder and full LDM architecture (inspired by rombach2022highresolution).
The LDM used in this study consists of a variational autoencoder (VAE) and a U-net (unet). The VAE performs the dimension reduction from the geomodel space to the latent space and vice-versa (Figure 2(a)). The U-net is used for the reverse diffusion process (Figure 2(b)). The VAE consists of an encoder and a decoder. The encoder is comprised of three downsampling convolutional layers, alternated with four residual blocks (group normalization – convolution – group normalization – convolution), such that the total downsampling ratio $f=8$ and the latent space $\bm{\xi}$ is of dimensions $N_{x}/8\times N_{y}/8$. Note that, in contrast to most VAE implementations, which rely on a 1D latent space, LDMs use a 2D latent space. The decoder mirrors the encoder, and thus has four residual blocks and three upsampling convolutional layers. The U-net consists of two downsampling residual blocks, one middle block with an attention mechanism (attention), and two upsampling residual blocks. The overall architecture follows the implementation of the original LDM in rombach2022highresolution. Our implementation is developed using the Python libraries diffusers and monai (see the ”Code availability” section for additional details). The detailed VAE and U-net architectures are shown in Figure 3.
@@ 2.4 LDM loss function and training
[…]
The latent variable $\bm{\xi}$ is intended to be standard normal (rombach2022highresolution), and the K-L divergence quantifies the difference between the distribution provided by the encoder and $\mathcal{N}(\mathbf{0},\mathbf{I})$, i.e.,
@@
[…]
In practice, during training, all losses are computed in batches of $N_{b}$ models. Pseudo-code for the U-net training is provided in Algorithm 1 (modified from ho2020denoising). For inference (generation), denoising starts from a lower dimensionality latent vector $\bm{\xi}_{T}$. The denoising in Eq. 7 is applied on $\bm{\xi}$. At the final step it is transformed by the decoder into a full-size geomodel $\mathbf{m}_{DM}$. This procedure is summarized in Algorithm 2 (modified from ho2020denoising).
@@
The dataset used in this study consists of 4000 conditional realizations, of dimensions $64\times 64$ ($N_{x}\times N_{y}=N_{c}=4096$), of three-facies geomodels. The facies correspond to high-permeability channel sand, low-permeability background mud, and intermediate-permeability levee features. The training realizations are generated with Petrel software (Schlumberger) using object-based modeling. These realizations are of similar style to those used by Song_Mukerji_Hou_2021. A 70-20-10% split for training, validation, and testing is applied. There are five hard-data locations, where the models are conditioned to channel facies. Five training realizations are shown in Figure 4.
The total training time is about 2 h (of which 0.5 h is for VAE training and 1.5 h is for U-net training) on one Nvidia A100 GPU. We use the Adam optimizer (adam), with a learning rate of 1e-4 and a batch size of 16. The scheduler for the noising process is linear, with $\beta_{1}=0.0001$ and $\beta_{T}=0.02$, as in ho2020denoising. A total of 100 DDIM steps are applied. The weights $\lambda_{KL}$ and $\lambda_{h}$ are set to $10^{-6}$ (as in rombach2022highresolution) and $10$, respectively.